OpenStudy (candy13106):

you are giving an expression for the volume of the rectangular prism. find an expression for the missing dimension

OpenStudy (candy13106):



jhonyy9 (jhonyy9):

let be the missing side Y you know that the volume of a rectangular prism is area of base time height so can you rewriting this knowledge using the missed side ?

jhonyy9 (jhonyy9):

area of base so area of a rectangle equal ?

OpenStudy (candy13106):

length times wihth

jhonyy9 (jhonyy9):

exactly so than you use what you know there note the missed side by one y

OpenStudy (candy13106):

ok so x+2 times x+4 times y?

jhonyy9 (jhonyy9):

just area of base equal y*(x+4) - yes ? and the volume will be equal with area of base time the height - yes

OpenStudy (candy13106):

ok

jhonyy9 (jhonyy9):

so than you can writing V = y(x+4)(x+2) - exactly how you have wrote above so for you can calcule the missed side just rewrite this equation for y - ok. ?

OpenStudy (candy13106):

k

OpenStudy (candy13106):

do i foil it

jhonyy9 (jhonyy9):

V=y(x+2)(x+4) divide both sides by (x+2)(x+4) and will get the equation for y

OpenStudy (candy13106):

wat

jhonyy9 (jhonyy9):

V y(x+2)(x+4) ----------- = ------------ yes ? but in the right part the (x+2)(x+4) from the top and (x+2)(x+4) (x+2)(x+4) bottom will simplified and will get V ---------- = y (x+2)(x+4) yes ?

OpenStudy (candy13106):

ok

jhonyy9 (jhonyy9):

do you understand it sure right ?

OpenStudy (candy13106):

yes

jhonyy9 (jhonyy9):

so than was my pleasure good luck bye bye